# ACDC: Automated Creation of Digital Cousins for Robust Policy Learning

### v2

@@ B.2 Real-to-Sim Scene Generation: Additional Results
Figure 9: Qualitative ACDC real-to-sim scene reconstruction results. Multiple cousins are shown with a robot collecting demonstrations. Images cropped by dashed squares are input RGB images.
Figure 10: Qualitative ACDC real-to-sim scene reconstruction results without ground truth camera intrinsics $\mathbf{K}$. Images cropped by dashed squares are input RGB images.
Figure 11: Qualitative ACDC real-to-sim scene reconstruction results without ground truth camera intrinsics $\mathbf{K}$. Images cropped by dashed squares are input RGB images.
@@
Additional qualitative results of our ACDC real-to-sim cousin creation and scene reconstruction pipeline are presented in Fig. 9. For multi-view visualizations, please refer to our accompanying video and website.
Our ACDC real-to-sim pipeline has the potential to create cousins and reconstruct scenes from a single RGB image without requiring ground truth camera intrinsics. We employ the Paramnet-360Cities-edina-uncentered model from PerceptiveFields [94] to estimate camera intrinsic matrix $\mathbf{K}$ from the input RGB image. Fig. 10 and Fig. 11 present the ACDC real-to-sim scene reconstruction results using the estimated $\mathbf{K}$. This capability may enable large-scale demonstration collection in the future by leveraging in-the-wild web images that lack ground truth camera intrinsics.
@@ B.3 Failure Cases
The first three limitations are directly tied to how ACDC is parameterized. For (a), because ACDC relies on relatively accurate depth estimations for computing predicted object 3D-bounding boxes, poorly estimated depth maps can result in correspondingly poor object model estimations from ACDC. Native depth sensors can struggle to produce accurate readings near object boundaries where discontinuities in the depth map may occur, and is compounded when an object has many fine boundaries, such as plants and fences. Moreover, because we rely on an off-the-shelf foundation model (DepthAnything-v2) to predict synthetic depth maps, we inherit its own set of limitations, such as poor predictions on esoteric objects or under adversarial visual conditions. Similar to (a), occlusion (b) becomes significant when it results in an inaccurate estimation of a given object’s overall bounding box. For some objects, such as cabinets and other furnitures, observing two faces is usually sufficient, but for other smooth objects, such as balls or plushes, occlusion can have nontrivial impacts on the corresponding generation from ACDC. Lastly, ACDC can struggle when there is a mismatch between object category labels from the input RGB image and the available object asset categories from our dataset. Because we do not enforce any naming or category-abstraction level from our dataset, our category-matching method (CLIP) may fail to associate categories due to esoteric naming schemes (e.g.: bottom_cabinet_no_top) or abstraction level mismatches (e.g.: cup vs. coffee_cup vs. drinking_cup vs. water_cup), resulting in suboptimal object asset candidates when selecting digital cousins.
@@ B.4 Comparison with URDFormer
  • URDFormer relies on accurate bounding box information which often requires manual human annotation, whereas ACDC is fully automatic and uses no human input.
In general, we find that while URDFormer can produce synthetic scene textures that visually match the real-world scene’s textures, ACDC can match or even outperform URDFormer’s ability to spatially reconstruct a given scene accurately, while additionally being object-agnostic (and thus able to detect and generate a much more diverse set of object categories) and fully automatic with no manual human annotation.
@@ B.6 Sim-to-Sim Policy Learning with Digital Cousins
An unexpected behavior is observed in the Drawer Opening task, where the 4-cousin policies perform sub-optimally. We believe this is due to the limited number of cabinets with drawers available for cousin selection. Among the four cousins, the first two are geometrically similar, as are the last two, but there is a significant similarity gap between the second and third cousins. This is partially illustrated by their DINO embedding distances to the digital twin: 7.78, 9.32, 14.10, and 14.90. The demonstrations collected on these four assets may not form a high-quality distribution for training. In contrast, the 4-cousin policy in the Door Opening task yield decent results, likely because there are more than 40 assets available for cousin selection, allowing ACDC to form a relatively narrower distribution. The geometric similarities between the four cousins in the Door Opening task are more continuous in terms of DINO similarity to the digital twin, with DINO distances being 6.49, 7.51, 8.13, and 9.66. However, 8-cousin policies still performed well in this relatively limited category, much better than all-assets policies and twin policies. A key takeaway is that: (1) when there are a sufficient number of assets to choose cousins from, all cousin policies can outperform twin policies on held-out cousins, and (2) more cousins should be found when the number of available assets is relatively small for the target category.

### v3

@@ B.2 Real-to-Sim Scene Generation: Additional Results
Figure 9: Qualitative real-to-sim digital cousin scene generation results. Multiple cousins are shown with a robot collecting demonstrations. Images cropped by dashed squares are input RGB images.
Figure 10: Qualitative real-to-sim digital cousin scene generation results without ground truth camera intrinsics $\mathbf{K}$. Images cropped by dashed squares are input RGB images.
Figure 11: Qualitative real-to-sim digital cousin scene generation results without ground truth camera intrinsics $\mathbf{K}$. Images cropped by dashed squares are input RGB images.
@@
Additional qualitative results of our real-to-sim digital cousin creation and scene generation pipeline are presented in Fig. 9. For multi-view visualizations, please refer to our accompanying video and website.
Our real-to-sim digital cousin creation pipeline has the potential to create cousins and reconstruct scenes from a single RGB image without requiring ground truth camera intrinsics. We employ the Paramnet-360Cities-edina-uncentered model from PerceptiveFields [94] to estimate camera intrinsic matrix $\mathbf{K}$ from the input RGB image. Fig. 10 and Fig. 11 present the ACDC real-to-sim digital cousin scene generation results using the estimated $\mathbf{K}$. This capability may enable large-scale demonstration collection in the future by leveraging in-the-wild web images that lack ground truth camera intrinsics.
@@ B.3 Failure Cases
The first three limitations are directly tied to how ACDC is parameterized. For (a), because ACDC relies on relatively accurate depth estimations for computing predicted object 3D-bounding boxes, poorly estimated depth maps can result in correspondingly poor object model estimations. Native depth sensors can struggle to produce accurate readings near object boundaries where discontinuities in the depth map may occur, and is compounded when an object has many fine boundaries, such as plants and fences. Moreover, because we rely on an off-the-shelf foundation model (DepthAnything-v2) to predict synthetic depth maps, we inherit its own set of limitations, such as poor predictions on esoteric objects or under adversarial visual conditions. Similar to (a), occlusion (b) becomes significant when it results in an inaccurate estimation of a given object’s overall bounding box. For some objects, such as cabinets and other furnitures, observing two faces is usually sufficient, but for other smooth objects, such as balls or plushes, occlusion can have nontrivial impacts on the corresponding generation of digital cousins. Lastly, ACDC can struggle when there is a mismatch between object category labels from the input RGB image and the available object asset categories from our dataset. Because we do not enforce any naming or category-abstraction level from our dataset, our category-matching method (CLIP) may fail to associate categories due to esoteric naming schemes (e.g.: bottom_cabinet_no_top) or abstraction level mismatches (e.g.: cup vs. coffee_cup vs. drinking_cup vs. water_cup), resulting in suboptimal object asset candidates when selecting digital cousins.
@@ B.4 Comparison with URDFormer
  • URDFormer relies on accurate bounding box information which often requires manual human annotation, whereas ACDC is fully automated and uses no human input.
In general, we find that while URDFormer can produce synthetic scene textures that visually match the real-world scene’s textures, ACDC can match or even outperform URDFormer’s ability to spatially reconstruct a given scene accurately, while additionally being object-agnostic (and thus able to detect and generate a much more diverse set of object categories) and fully automated with no manual human annotation.
@@ B.6 Sim-to-Sim Policy Learning with Digital Cousins
An unexpected behavior is observed in the Drawer Opening task, where the 4-cousin policies perform sub-optimally. We believe this is due to the limited number of cabinets with drawers available for cousin selection. Among the four cousins, the first two are geometrically similar, as are the last two, but there is a significant similarity gap between the second and third cousins. This is partially illustrated by their DINO embedding distances to the digital twin: 7.78, 9.32, 14.10, and 14.90. The demonstrations collected on these four assets may not form a high-quality distribution for training. In contrast, the 4-cousin policy in the Door Opening task yield decent results, likely because there are more than 40 assets available for cousin selection, allowing reconstructed digital cousins to form a relatively narrower distribution. The geometric similarities between the four cousins in the Door Opening task are more continuous in terms of DINO similarity to the digital twin, with DINO distances being 6.49, 7.51, 8.13, and 9.66. However, 8-cousin policies still performed well in this relatively limited category, much better than all-assets policies and twin policies. A key takeaway is that: (1) when there are a sufficient number of assets to choose cousins from, all cousin policies can outperform twin policies on held-out cousins, and (2) more cousins should be found when the number of available assets is relatively small for the target category.